OpenStudy (anonymous):

If the point (x,square root 3/3) is on the unit circle, what is x?

terenzreignz (terenzreignz):

A point (x , y) is on the unit circle if and only if it satisfies this relation : x² + y² = 1 Now, your y is given to be sqrt(3) / 3 So plug that in, and find x.

OpenStudy (anonymous):

the choices are a:2/3 b:square root(6)/3 c:4/9 d:square root(2)/3

terenzreignz (terenzreignz):

So... which did you get? Plug in sqrt(3)/3 for y, and then solve for x.

OpenStudy (anonymous):

i cant figure it out can u do it for me please

terenzreignz (terenzreignz):

\[\huge x^2+y^2=1\] Now, we have \[\large y=\frac{\sqrt{3}}{3}\]and so we plug that in \[\huge x^2+\left(\frac{\sqrt{3}}{3}\right)^2=1\] Giving us.... \[\huge x^2+\frac{3}{9}=x^2+\frac{1}{3}=1\]

OpenStudy (anonymous):

so which choice is right

terenzreignz (terenzreignz):

Just solve for x. You can do it :)

OpenStudy (anonymous):

is the answer 2/3

OpenStudy (anonymous):

u there

terenzreignz (terenzreignz):

Yeah... thinking.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

think faster i have so olong to do this

terenzreignz (terenzreignz):

Okay, the answer is in the choices all right. I was having doubts.

OpenStudy (anonymous):

is the answer 2/3

terenzreignz (terenzreignz):

I'll believe you if you can show how you got it :)

OpenStudy (anonymous):

2/3^2+3/9=1

terenzreignz (terenzreignz):

Nu-uh... (2/3)² + (3/9) = (4/9) + (3/9) = 7/9 which is not equal to 1. Try again.

OpenStudy (anonymous):

how bout 4/9^2+3/9=1

terenzreignz (terenzreignz):

It looks like you're guessing. Solve for x from the equation \[\huge x^2 + \frac{1}{3}=1\] And in time, you'll be able to answer questions like this even without choices.

OpenStudy (anonymous):

no im not so is that wrong i dont understand this

OpenStudy (anonymous):

yololol do u know the answer

terenzreignz (terenzreignz):

Find out. Work out \[\huge \left(\frac{4}{9}\right)^2 + \frac{1}{3}\] And see if it equals 1. Yes, I do know the answer, but I'd prefer, and it'd be better for you, if you could arrive at the answer yourself, if even with a little bit of help from me :)

OpenStudy (anonymous):

please if u tell me the answer ill try to see how u got it

OpenStudy (anonymous):

the answer is d

terenzreignz (terenzreignz):

Sorry, the power went out... Have a little more faith in yourself... YOU CAN DO THIS :)

OpenStudy (anonymous):

so thats wrong

OpenStudy (anonymous):

then it is b that was my first choice i thought

OpenStudy (anonymous):

i figure it out 6/9+3/9=1

OpenStudy (anonymous):

so the answer is b

OpenStudy (anonymous):

thsanks for the help

terenzreignz (terenzreignz):

See, you can do it :)

OpenStudy (anonymous):

thanks for believing in me im going to close this now